OpenStudy (bambimonster):

A large lecture theater has 230 seats, 23 of which can accomodate left-handed students. Suppose it is known that 13% of people are left-handed. One class held in the theater has 209 students. (a) Let X be the number of left-handed students in the class. The mean and standard deviation of X are, respectively,

OpenStudy (kropot72):

The probability of a student being left-handed is 0.13. The binomial distribution applies to this question. The binomial distribution can be approximated by the normal distribution in this case where \[mean=np=209\times0.13\] and \[standard\ deviation=\sqrt{np(1-p)}=\sqrt{209\times0.13\times0.87}\]

OpenStudy (kropot72):

Note: It seems that any overflow of the seats for left-handed students would need to be accommodated in the seats for the right-handed.

OpenStudy (anonymous):

@kropot72 and that is likely to happen with a high probability (77% using exact binomial for P(x>=24)) All that besides all the guys that never take notes in class and do not really care whether the chair is for left or right-handed :)

OpenStudy (kropot72):

@CarlosGP Agreed :0

OpenStudy (bambimonster):

i got it already thanks a bunch

OpenStudy (kropot72):

You're welcome :)